OpenStudy (anonymous):

How do I express this equation into simplest radical form? χ÷√3=√6? The answer is 3√2 but why and how?

OpenStudy (anonymous):

isolate x by multiplying \[\sqrt{3} * \sqrt{6}\]

OpenStudy (anonymous):

\[\sqrt{a}*\sqrt{b}=\sqrt{a*b}\]

OpenStudy (anonymous):

ohh.. wait so I just transfer sq. root of 3 to the right? There's nothing i have to do with the x/sq.root of 3?

OpenStudy (anonymous):

You multiply both sides by \[\sqrt{3}\]

OpenStudy (anonymous):

This will isolate x, then you can simplify the square root of 18

OpenStudy (anonymous):

ahh, i see. thank you!

OpenStudy (phi):

notice they want simplest radical form so \[\frac{x}{\cancel{\sqrt{3}}}\cdot \cancel{\sqrt{3}}= \sqrt{6\cdot3}= \sqrt{2\cdot3\cdot3}= 3\sqrt{2}\]